OpenStudy (anonymous):

A 720 kg race car starts from rest on the race track and travels 39 m in 2.9 seconds. The car is uniformly accelerated during the entire time. (a) What is the acceleration of the car? (b) What is the net force acting on the car? (c) What is the agent of this force, i.e., what object pushes the car forward? (a single-word noun)

OpenStudy (alfie):

\[x_f = x_i +v_i t+\frac{1}{2}at^2\] There is no initial velocity. And we suppose the initial position is 0 . So we have. \[a = \frac{2x_f}{t^2}\] So a = 9,27m/s^2. There is no friction, so the only force is given by the acceleration of the car or: F = ma. => F = 720*9.27 = 6677N

OpenStudy (anonymous):

x=.5at^2+vt 3.9=.5*(2.9)^2a = a= 9.27 m/s^2 f=ma f=9.27*720=6674.4N that force cause to move car .